HUBBLE SPACE TELESCOPE DAILY REPORT #2601 PERIOD COVERED: 0000Z (UTC) 04/12/00 - 0000Z (UTC) 04/13/00 Daily Status Report as of 104/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-1. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. As described in 2.1 and HSTAR 7575, the acquisition for the final iteration of this proposal failed, affecting three observations. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.3 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.4 Completed Four Sets of WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.5 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.6 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.7 Completed STIS/CCD 8284 (Deep STIS Observations of BALQSO PG 0946+301) The Space Telescope Imaging Spectrograph (CCD ) was used to study AGN outflows to determine of the ionization equilibrium and abundances (IEA) in broad absorption line {BAL} QSOs. In the last few years it has become clear that partial coverage of the continuum source and saturation are common in BALs, and these frustrate attempts to determine the IEA from available HST data because of limitations imposed by the S/N ratio and spectral coverage. We will re-observe the most suitable BAL IEA target: PG 0946+301. The observation completed with no reported problems. 1.8 Completed Five Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD/MA1 8424 (MAMA Sensitivity and Focus Monitor C8) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of object GRW+70D582 in order to measure the sensitivity and to focus monitor C8. This also is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes, and to monitor the STIS focus in a spectroscopic mode. A wavelength calibration was also taken. The observations were completed as planned, and no anomalies were reported. 1.10 Completed WF/PC-2 8390 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was to be used to make observations of the planetary nebula NGC7009 in order to study its parallax distance. As described in 2.1 and HSTAR 7575, the acquisition for this proposal failed, affecting nine observations. 1.11 Completed STIS/CCD/MA2 7359 (Chemical Abundances in QSO Broad Absorption Line Regions) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the quasar QSO-143648 in order to study the chemical abundances in the QSO's broad absorption line regions. These observations are designed to test the possible models for high metallicity in BALQSOs and, indirectly, test the origins of the BAL phenomenon. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.12 Completed STIS/MA2 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 8448 (Cycle 8 INTFLAT Sweeps, VISFLAT Mini-Sweep, and Linearity Test) WFPC2 was used to monitor the pixel to pixel flatfield response and the VISFLAT lamp degradation, as well as detect any possible changes due to contamination. The observations completed nominally. 1.14 Completed STIS/CCD 8168 (The Optical Counterpart of the Soft Gamma-ray Repeater 1900+14) The Space Telescope Imaging Spectrograph (CCD) was used to observe soft gamma-ray repeater {SGR} 1900+14, one of only four known, which are likely an exotic class of neutron star known as magnetars with ultra-strong surface magnetic fields. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 6 HSTAR 7575 documents a guide star acquisition failure at 103/122326Z. This occurred when the scan step limit was exceeded on FGS-2. The proposals described in 1.2 and 1.10 were affected. Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 17 Successful: 17 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared once. There was a STIS EMC retry at 103/1215542Z. The STIS flight software error counter was cleared at 103/1237Z per ROP NS-12. A TTR was written for a negative mode 1 acquisition at 104/011400Z. Per an operations note steps were taken to successfully re-acquire data. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW